### v1

# It’s complicated. The relationship of algorithmic fairness and non-discrimination regulations in the EU AI Act

Abstract.
What constitutes a fair decision? This question is not only difficult for humans but becomes more challenging when Artificial Intelligence (AI) models are used. In light of discriminatory algorithmic behaviors, the EU has recently passed the AI Act, which mandates specific rules for AI models, incorporating both traditional legal non-discrimination regulations and machine learning based algorithmic fairness concepts. This paper aims to bridge these two different concepts in the AI Act through: First a high-level introduction of both concepts targeting legal and computer science-oriented scholars, and second an in-depth analysis of the AI Act’s relationship between legal non-discrimination regulations and algorithmic fairness. Our analysis reveals three key findings: (1.), most non-discrimination regulations target only high-risk AI systems. (2.), the regulation of high-risk systems encompasses both data input requirements and output monitoring, though these regulations are often inconsistent and raise questions of computational feasibility. (3.) Regulations for General Purpose AI Models, such as Large Language Models that are not simultaneously classified as high-risk systems, currently lack specificity compared to other regulations. Based on these findings, we recommend developing more specific auditing and testing methodologies for AI systems. This paper aims to serve as a foundation for future interdisciplinary collaboration between legal scholars and computer science-oriented machine learning researchers studying discrimination in AI systems.
EU AI Act, Algorithmic Fairness, Discrimination, Bias, Fundamental Rights
1. Introduction
How can we ensure fair algorithms in the context of AI Systems? When regulating algorithms and, more specifically, Artificial Intelligence (AI) systems, this question becomes fundamental. While the question is also crucial in an analog world, it becomes increasingly pressing in the digital world. The European Union (EU) has become one of the forerunners of regulations in the digital age with regulatory frameworks, such as the Digital Service Act [^1] Regulation (EU) 2022/2065., the Digital Markets Act [^2] Regulation (EU) 2022/1925., or the Data Act [^3] Regulation (EU) 2023/2854.. In order to maintain fair data processing in the machine learning domain, the European Union (EU) recently adopted the Artificial Intelligence Act (AIA) [^4] Regulation (EU) 2024/1689.. The aim of this AI regulation is to maintain a level playing field for “ethical principles” (Nolte et al., 2024) within and outside the European Union (see recital 27 AIA).
Fair algorithmic processing matters also from a computational perspective. Over the past decade, algorithmic fairness has become a well-established field within machine learning that focuses on defining, mitigating, and evaluating discriminatory behavior of Artificial Intelligence models (Pessach and Shmueli, 2023). The importance becomes clear when looking at several discriminatory behaviors of algorithms in the past, ranging from hiring (Dastin, 2022) to Social Welfare systems (Hadwick and Lan, 2021). In recent years, large generative (multimodal) models, particularly Large Language Models (LLMs) with their high accessibility and wide range of applications have posed significant new challenges to the algorithmic fairness domain (Chu et al., 2024; Kotek et al., 2023).
The relationship between computer science and legislation for efficient regulations within the AI Act is complex. While the interplay between algorithmic fairness and non-discrimination law is often unclear (Weerts et al., 2023), efficient AI legislation requires interdisciplinary collaboration connecting the legal domain and computer science. Although, legal and technical concerns regarding algorithmic fairness and non-discrimination law have a long history (Barocas and Selbst, 2016; Hacker, 2018; Wachter et al., 2021), recent developments have accelerated the pace. An increasing number of companies and laypersons are using these technologies, and with increased usage, the likelihood of misuse also rises. Thus, the need for efficient and effective regulations to protect individuals is substantial.
Previous literature: Many papers on discrimination and AI, limited amount on the interaction of the AI Act and discrimination. Discrimination in algorithms is not a novel phenomenon and was analyzed in a pre-AI Act era. Already in the 1980s, algorithmic discrimination was observed in admission settings (Connors et al., 1981; Williams et al., 1981), and gender inequalities in educational software were discussed (Huff and Cooper, 1987). In 1996, researchers pointed out the levels at which such technical constraints and social institutions’ biases can occur (Friedman and Nissenbaum, 1996). More recently, the interaction between AI and automated decision-making systems with traditional European and US non-discrimination law has been studied extensively. Barocas and Selbst (2016) presented a taxonomy of different sources of discrimination and their impact on humans. The relationship among various European non-discrimination laws has also been studied extensively (Wachter et al., 2021, 2020; Hacker, 2018; Weerts et al., 2023; Xenidis and Senden, 2020). For example, the seminal study by Wachter et al. (2021) presented a fairness metric that connects algorithmic fairness to the jurisprudence of the European Court of Justice in relation to the General Data Protection Regulation. Similarly, the connection between non-discrimination law and the GDPR has been analyzed to “unlock the algorithmic black box” (Hacker, 2018). Naturally, since these papers were published before the AI Act was passed, they do not contain references to it.
Recently, several papers have already addressed the AI Act. Short(er) studies such as (Deck et al., 2024a; Ruohonen, 2024) have analyzed the relationship between fairness and the AI Act. Additionally, sections within papers of broader scope beyond pure non-discrimination analyses have mentioned aspects of the relationship between algorithmic fairness and the AI Act (Wachter, 2024; Hacker et al., 2024b; Novelli et al., 2024) . The AI Act and its relation to gender equality and non-discrimination law have also been discussed (Lütz, 2024). Most similar to our paper in terms of the AI Act, Bosoer et al. (2023) investigated the non-discrimination regulations in the draft of the AI Act without a strong focus on the interaction with computer science. However, their focus was on the draft, whereas our paper focuses on the substantially different final version. Most similar in terms of the interaction between computer science and legal research is the paper of Weerts et al. (2023). In contrast to our work they did not focus on the AI Act.
Our contributions: From regulatory clarity to vagueness This paper aims to foster and extend an interdisciplinary view on explicit and implicit fairness regulations in the AI Act regarding algorithmic fairness in the ML domain. We aim to introduce the field of algorithmic fairness to legally oriented researchers in Section 2.1. Afterwards, we introduce non-discrimination law to computer scientists in Section 2.2 and discuss the truncated relationship between the two fields in Section 2.3.
For the in-depth analysis of the EU AI Act non-discrimination regulations in Section 3, our contributions are as follows (Figure 1 shows our results and findings in a nutshell):
  1. (1)
    We present the history, scope, and intentions of non-discrimination regulations in the AI Act (Sec. 3.1).
  2. (2)
    We identify which issues of algorithmic fairness are addressed explicitly and implicitly within the AI Act (Sec. 3.2).
  3. (3)
    We find that high-risk AI Systems (Article 6 AIA) are most extensively regulated regarding non-discrimination. Regulations are partly explicit and partly implicit. Overall the regulations are inconsistent with respect to algorithmic fairness, with regulations on input and output data at different stages (Sec. 3.3).
  4. (4)
    Non-discrimination regulations of General Purpose AI models (GPAI) in Article 53-56 AIA remain insufficiently specific and lack clear definition (Sec. 3.4).
We summarize our findings and give recommendations for future research in Section 4. Additionally, based on our results, we propose approaches to overcome some of the challenges and sketch possible actions to address the inconsistencies. This includes a call for closer interactions between computer scientists and legal researchers.
Figure 1. Results of our analysis of the EU AI Act’s non-discrimination regulation and algorithmic fairness in a nutshell: Low/medium risk systems [^5] The AI Act also includes a third category of AI Models in Article 51 AIA: certain AI Systems. However, certain AI Systems are not target of any specific non-discrimination regulation and thus not included in our analysis.  are not regulated in terms of non-discrimination obligations within the AI Act. Key terms for high-risk systems are the partly well-defined concept of non-discrimination and bias (less well defined). High-risk systems, as well as GPAI models, are also regulated with respect to Fundamental Rights. GPAI models, with the key term systemic risk, show the lowest regulatory clarity. Overall, regulatory clarity for non-discrimination matters decreases from high-risk systems to GPAI models.
2. A primer on non-discrimination from an AI and a legal perspective
Interdisciplinary research on algorithmic fairness is challenging, both from a computer science and a non-discrimination law perspective. Understanding legal reasoning can be challenging without prior legal knowledge, just as understanding algorithmic methods is difficult without a computational background.
However, the complexity in interdisciplinary work extends beyond pure knowledge. The challenges begin with terminology: For instance, in computer science, “fairness” is a term that can refer to different desiderata that aim to prevent socially or morally undesirable behavior or outcomes of algorithms. Although fairness has been discussed as principle in economic law contexts (Scheuerer, 2023), it is not a specific legal term. The arguably closest legal counterpart term is “non-discrimination”, which focuses on preventing unfair treatment based on characteristics such as race, gender, or other attributes on legal grounds. The interaction between both key concepts from law and computer science is demanding.
Thus, we highlight the differences and common grounds of algorithmic fairness and non-discrimination regulations in the following sections. After presenting an introduction to algorithmic fairness in Section 2.1 — for legal researchers — and an introduction to EU non-discrimination law in Section 2.2 — for computer science oriented researchers —, we try to bridge the gap between these two areas in Section 2.3.
2.1. Algorithmic fairness: A gentle introduction
Algorithmic Fairness has emerged as an established computer science and AI related field over the past years. Algorithmic fairness focuses on uncovering and rectifying disadvantageous treatment of individuals or groups in machine learning models (Mitchell et al., 2021). This treatment must be considered within appropriate social, theoretical, and legal contexts (Pessach and Shmueli, 2023) and includes evaluation and auditing methods to test for unfairness.
Origins of unfairness in algorithms. The origins of unfairness are manifold. Unfair algorithmic behavior can have different sources. In order to discuss the origins, another — not well defined — term from computer science is often used: biases. There exist many different bias definitions. Since the EU AI Act does not define it itself, we use a definition from the EU commission from 2021. According to this, a bias can be defined as “[…] bias describes systematic and repeatable errors in a computer system that create unfair outcomes, such as favouring one arbitrary group of users over others” (Commission, 2021). Biases are primarily discussed within data used for AI models and can be sorted into different categories. Most researchers emphasize the different aspects of data inequalities when discussing algorithmic fairness and sort data biases into different categories, such as the over-representation of specific groups in datasets (Zou and Schiebinger, 2018). Hacker (2018) sorts the biases in data into two groups: biased training data vs. unequal ground truth; others use more categories. Barocas and Selbst (2016) uses five categories to map biases, while Mehrabi et al. (2021) uses 21 different categories for biases. It is important to note that the data are only — one albeit important — source of biases. For example, design decisions in the algorithms or structural power asymmetries also lead to algorithmic unfairness (Mehrabi et al., 2021; Gebru et al., 2021; Sousa e Silva, 2024).
Approaches to algorithmic fairness: Quantification. Based on the different understandings of unfairness and for the purpose of uncovering and rectifying unjustified treatment between groups in algorithms, so-called fairness metrics have been proposed (Corbett-Davies et al., 2023; Castelnovo et al., 2022; Verma and Rubin, 2018). The core idea of fairness metrics is to quantify algorithmic outputs and make them comparable through numerical measurements. Most metrics try to assess whether an output of an ML system is unfair to individuals (Dwork et al., 2012) or groups of people (Barocas et al., 2023; Binns, 2020). They provide a quantitative measure intended to indicate whether an algorithm shows unjustified unequal treatment. Fairness metrics also include moral norms (Deck et al., 2024a; Hellman, 2020).
The emergence of the field of algorithmic fairness has challenged existing approaches to ML by highlighting the need to contextualize them, decide on and reason for the optimization criteria chosen. First, it has been shown that an apparent objective number needs to be contextualized to the application under consideration (Wachter et al., 2021). Furthermore, some metrics are, from a mathematical viewpoint incompatible with each other [^6] Please note that they are also incompatible from a moral point of view (Heidari et al., 2019)., thus some fairness metrics cannot be fulfilled at the same time (Chouldechova, 2017; Kleinberg et al., 2016). This can lead to an effect of “d-hacking” or “fairness hacking” where users can choose their favorite metric to create the (technical and mathematical) impression that the algorithm is fair (Black et al., 2024; Meding and Hagendorff, 2024). Fairness metrics for mitigation can play a role at different stages of the development, testing and deployment of an algorithm. The pre-processing stage, the in-training stage, and the post-processing phases are differentiated (Barocas et al., 2023; Binns, 2018). In the pre-processing phase, the input data itself is altered to ensure fair processing (Caton and Haas, 2024; Kamiran and Calders, 2012). For the in-training phase, the optimization process during the training of an ML model is adjusted to include fairness as an optimization goal. Finally, in the post-processing phase, the output of an already trained model is adapted, see Figure 2 in the appendix. The first and the last approach makes it possible to perform fairness analyzes even if one only has black box access to the model (Caton and Haas, 2024).
The era of LLMs further complicates the situation. LLMs have the advantage — or disadvantage depending on the viewpoint — that their exact use case is most of the time not predefined. They are trained on large amounts of various data. This diversity makes them later applicable to different contexts, from providing cookie recipes to coding exercises. However, the diversity then also includes variations in unfairness. Algorithmic fairness in relation to LLMs is thus somewhat different from algorithmic fairness in classical AI (Doan et al., 2024; Chu et al., 2024). Some researchers have applied classical fairness metrics in classification settings to LLMs (Chhikara et al., 2024). Additionally, it was shown that LLM-specific issues arise due to the use of massive data and processing (Kotek et al., 2023; Ferrara, 2023; Navigli et al., 2023; Huang et al., 2024) at various steps of algorithmic development. Thus, it was argued that LLMs cannot yield fair outcomes at all (Anthis et al., 2024).
2.2. EU non-discrimination Law: A gentle introduction, too.
Law is different from computer science. Nevertheless — of course — legal researchers have also discussed different concepts of discrimination. In contrast to the term “fairness”, the term “(non-)discrimination” is a specific legal concept  [^7] For an in-depth introduction we refer to the paper of Weerts et al. (2023) or Wachter et al. (2020).. For a better understanding of these legal concepts, we first describe traditional EU non-discrimination law as well as the Fundamental Rights approach of the EU. Both influence the AI Act.
Direct and indirect discrimination in EU Law. EU Law has a long history of non-discrimination regulations and many different laws include non-discrimination regulations [^8] See for an overview: https://commission.europa.eu/aid-development-cooperation-fundamental-rights/your-fundamental-rights-eu/know-your-rights/equality/non-discrimination_en. Primary EU law already includes non-discrimination regulations, one example is the prohibition of unequal payment based on gender (Article 157(1) TFEU). In order to legally assess discrimination under EU Law, traditionally two different versions of discrimination are differentiated: direct and indirect discrimination (Zuiderveen Borgesius et al., 2024; Wachter et al., 2020). In context of the US law, these two concepts correspond to disparate treatment (direct discrimination) and disparate impact (indirect discrimination) (Weerts et al., 2023). However, importantly, EU-based direct discrimination does not need — compared to US-based disparate treatment — any intentional wrongdoing (Weerts et al., 2023; Xenidis and Senden, 2020; Adams-Prassl et al., 2023).
The AI Act is based on the legal approach of discrimination-related regulations (Deck et al., 2024a). Thus, when we analyze the EU AI Act regulations in this paper, we will minimize the usage of the vague CS related concept of fairness and instead focus on the legally related term of non-discrimination.
Direct discrimination is defined as situations in which “one person is treated less favourably than another is, has been or would be treated in a comparable situation” (Article 2(2)(a) Directive 2000/43/EC). This means that individuals are discriminated against (directly) by using a (protected) attribute. In algorithmic contexts, an example would be a simple algorithm that grants anyone below 50 years a loan and denies it to anyone above 50 years. In this setting, only the protected attribute age is used to define the loan threshold. People would be directly discriminated against based on their age.
Indirect discrimination is much more challenging to tackle. In indirect discrimination cases, seemingly neutral attributes are used, but they rely on a protected attribute. An example of indirect discrimination would be if not the age is used but the amount of gray hair, which is usually highly correlated with age. If an algorithm does not use age directly but the amount of gray hair, then people are indirectly discriminated against based on their age.
Also, a third category of discrimination needs to be taken into account which challenges traditional non-discrimination law: intersectional discrimination. Intersectional discrimination concerns cases “originating in several inextricably linked vectors of disadvantage” (Xenidis, 2023). This becomes especially important since (modern) AI models do not use single variables as input — such as our simplified loan example based on the single variable above — but instead use many different aspects as input. This can lead to effects where discrimination may occur at the intersection of gender and age. Importantly, if you only investigate age and gender separately, no discrimination would be recognized (Weerts et al., 2023). It is an open question if and how, at the current stage, the European Court of Justice recognizes intersectional discrimination as a form of discrimination (Weerts et al., 2023; Xenidis, 2023; Atrey, 2018).
With these three types of discrimination based on EU discrimination law, the last concept that needs to be discussed is the concept of Fundamental Rights which is mentioned in multiple articles in the AI Act.
Fundamental Rights in EU law. The Charter of Fundamental Rights of the European Union (CFR) is the main Fundamental Rights regulation within EU law . Through Article 6(1) of the Treaty on European Union, the Charter is part of the primary law of the EU.
The CFR also includes non-discrimination law. For non-discrimination setting obligations, Article 21 CFR states: “Any discrimination based on any ground such as sex, race, color, ethnic or social origin, genetic features, language, religion or belief, political or any other opinion, membership of a national minority, property, birth, disability, age or sexual orientation shall be prohibited.”
However, the right to non-discrimination is not absolute. It is important to note that the right to non-discrimination is only one right in the CFR, among others, such as the freedom of expression and information (Article 11(1) CFR). In order to apply the CFR, rights need to be balanced (Article 52(1) CFR). First, legislators must balance different CFR rights proportionally when creating new laws. Additionally, when laws reference the CFR, implementers must also follow this balancing obligation. A logic of proportionality of the different rights guides Fundamental Rights assessments (Almada and Petit, 2023). The Fundamental Rights are not absolute, hierarchical, or quantifiable and need to be applied case by case (Sousa e Silva, 2024). This balancing and proportionality can make it very difficult to grasp the exact concept of Fundamental Rights.
2.3. Interaction of law and computer science: Differentiating between concepts.
Law and computer science are different fields. While this becomes clear from the different viewpoints above, it also leads to severe challenges when discussing algorithmic fairness and law: the terminology differs. The difficulty stems from the fact that different fields and society at certain times and places define fairness and discrimination in different ways (Sousa e Silva, 2024). While the compatibility of fairness metrics with normative legal statements has already been discussed (Deck et al., 2024a; Hellman, 2020; Wachter et al., 2021, 2020), there is still an inherent difficulty in combining these two fields. This interdisciplinary challenge needs to acknowledge that computer science fairness and legal non-discrimination regulation are moving in the same direction; however, they present different concepts (Deck et al., 2024b). Although inspired by law, the technical algorithmic fairness debate primarily focuses on detecting, preventing and mitigation unfairness in algorithms(Deck et al., 2024b), while legal non-discrimination law focuses on normative statements.
As shown in the previous section, the algorithmic fairness domain uses terms like bias or fairness. These concepts are difficult to grasp and are topics of debate regarding their exact meaning (Sousa e Silva, 2024; Genovesi et al., 2024; Xenidis and Senden, 2020; Bosoer et al., 2023). Furthermore, in legal terms, algorithmic fairness is rarely used.
When we discuss algorithmic fairness and non-discrimination regulation, we are going to differentiate terms. First, (algorithmic) fairness refers to the AI-related domain of analyzing, evaluating, auditing, and mitigating unequal treatment between (groups of) individuals based on normative notions (Laux et al., 2024; Estévez Almenzar et al., 2022). When we use the terms (non-)discrimination, we will refer to classical non-discrimination law in the European Union.
3. EU AI Act’s non-discrimination regulations.
3.1. The EU AI Act: History, Scope, and Definitions
Emergent technology needs efficient regulation. Before the work on the European AI Act started, the EU High-Level Expert Group on Artificial Intelligence had already developed ethical guidelines for AI systems. These guidelines were later further developed for the AI Act (Koumeri et al., 2023). While the AI Act is not the first AI regulation worldwide (Wodi, 2024), it presents harmonized rules and obligations for the use and “placing on the market” (Article 1(2) AIA) of AI systems in the European Union. Thus, a level playing field for AI products within the Union is established. The AI Act is (predominantly) a product safety law (Almada and Petit, 2023). In fact, the original drafts did not have any individual rights components (Hacker et al., 2024b). The Members of the European Parliament did not prioritize non-discrimination regulation in the beginning (Chiappetta, 2023).
Risk is a central concept within the AI Act (Hacker et al., 2024a). Similar to other digital regulations in the EU, the AI Act defines risk in Article 3(2) AIA. A risk “means the combination of the probability of an occurrence of harm and the severity of that harm”. Based on this definition, the AI Act differentiates between different risk types with special requirements, which is not uncontroversial (Bosoer et al., 2023): prohibited systems (Article 5 AIA), high-risk systems (Article 6 ff. AIA), certain AI systems (Article 50 AIA), and GPAI models (Article 51 ff. AIA). Prohibited systems pose an unacceptable risk (Nolte et al., 2024). In contrast, high-risk systems pose a risk that can be handled through extensive regulations. Certain AI Systems are systems that can imitate human content, are emotion recognition systems, or biometric categorization systems, and therefore users must be informed transparently about their use. GPAI Models are, in computer science terms, (Multimodal) Foundation models, which are a separate category due to their capability to serve various purposes, both for direct use and integration in other AI systems (Article 3(66) AIA). If a system does not fall into one of these categories, only a limited set of rules applies. One example is the obligation that providers and deployers of AI systems ensure the AI literacy of their staff or other persons dealing with the system (Article 4 AIA).
The AI Act has specific wording for the difference between model and system. On the one hand, the AI Act defines in Article 3(1) AIA an AI system as: “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments”. While the AIA sticks to the AI system definition when introducing high-risk systems, it refers to GPAI models in Article 51 ff. AIA. For GPAI Model Rec. 97 AIA defines that “Although AI models are essential components of AI systems, they do not constitute AI systems on their own. AI models require the addition of further components, such as for example a user interface, to become AI systems.” From a computer science perspective, it is unusual not to use the term “model” or “machine learning model”, which are the terms most often used in technical papers. Since we are discussing legal norms, we stick to definitions provided by the AI Act, see also (Nolte et al., 2024).
3.2. Non-discrimination regulations within the AI Act: From clarity to vagueness.
Analyzing non-discrimination regulations starts with identifying regulations. In total, we scanned the AI Act for non-discrimination-related terms (from specific to unspecific concepts): discrimination, Fundamental Rights, fairness, and bias. We noticed that within the definitions of Article 3 AIA, the serious incident (Article 3(49) AIA) and systemic risk (Article 3(65) AIA) refer to Fundamental Rights. Thus, these were included in our analysis as well.
While analyzing the regulations, we observed that terms appear in the majority of cases in the recitals and only very limited in the articles of the AI Act. The recitals are only considered helpful when interpreting the articles (Klimas and Vaiciukaite, 2008) but do not have legal binding force (noa, 2014). Therefore, our analysis does not focus on the recitals.
Table 1 presents our main results. Not all articles directly relate to specific AI regulations but sometimes regulate market authorities, for example, Article 28 AIA. Thus, we focused on core AI regulation and present the list of articles mentioning non-discrimination-related terms in Table 1. A full table with all articles of the AI Act including non-discrimination-related terms, can be found in the appendix in Table 2.
Article
Discrimination /
discriminatory
Fundamental
Rights
Fair /
fairness
Bias
Systemic
risk
Serious
incident
Specific Topic
General Introduction
1 Subject matter of AI Act
2 Scope of the AI Act
3 Definitions
High-risk Systems and its obligations
6 Classification rules for high-risk AI systems
7 Amendments to Annex III AI Act
9 Risk management system for High-Risk AI Systems
10 Data and data governance for High-Risk AI Systems
13 Transparency and provision of information
14 Human oversight of High-risk Systems
15 Accuracy, robustness and cybersecurity
17 Quality management system
26 Obligations of deployers of high-risk AI systems
27 Fundamental Rights impact assessment for high-risk AI systems
GPAI Models
51 Classification of AI models with systemic risk
52 Notification Procedure for GPAI models
53 Obligations for providers of general-purpose AI models
54 Authorised representatives of providers of GPAI models
55 Obligations of providers of GPAI Models with systemic risk
56 Codes of practice for GPAI Models
Table 1. Non-discrimination and algorithmic fairness-related terms in the AI Act for selected most important articles. A full overview is provided in 2 in the appendix. From left to right, regulatory clarity decreases.
This analysis shows, first, that the majority of non-discrimination regulations in the EU AI Act concern the regulation of high-risk systems. GPAI models are only regulated implicitly by the systemic risk and serious incident terms. This will have an impact on our further analysis: we focus on high-risk systems compared to GPAI models.
Second, most non-discrimination regulations are only implicitly addressed by the reference to Fundamental Rights. Discrimination and bias as threats are explicitly addressed only for high-risk systems. But here too, we find a relation to Fundamental Rights.
Third, the term “fairness” from algorithmic fairness is not used at all within the AI Act. On the one hand, this result is surprising. The fairness aspect reveals a gap between wording in the computer science and ML community and the legal field. On the other hand, however, as the non-discrimination rules are based on established rules within EU legislative frameworks, it is not a surprising finding.
After analyzing which parts of the AI Act contain non-discrimination regulations, we will now analyze the details for high-risk systems and general purpose AI systems in detail.
3.3. High-risk AI non-discrimination regulations.
What is a high-risk AI system? Article 6 AIA defines different aspects of high risks. In combination with Annex III, high-risk systems are, for example, intended to be used as safety components in the management and operation of critical digital infrastructure (Annex III(2) AIA) or systems intended to be used for the recruitment or selection of natural persons (Annex III(4)(a) AIA). The reasoning behind the specific regulations in Article 6 AIA is that systems which pose a specific risk to fundamental rights or safety need tighter regulation (Fraser and y Villarino, 2024) than other less risky systems. It is expected that 5%-15% [^9] For a critical assessment of these numbers see (Almada and Petit, 2023). of all AI systems fall under the category of high-risk systems (Commission, 2021).
We identified Articles 9, 10, and 15 AIA as main non-discrimination provisions of high-risk systems. These will be discussed in detail in the following paragraphs [^10] Please note that we do not differentiate here between the obligations for providers and those obligations for the deployers of AI systems, since we focus on the interaction of computer science and law.. All other Articles listed in tab 1 will be discussed and summarized in the last paragraph.
Article 9 AIA: Undefined discrimination regulation within risk management systems Article 9 of the AI Act covers risk management systems of High-Risk Systems. According to Article 9(2)(a-d) AIA, risk management systems are designed to identify, evaluate, and mitigate risks of AI systems in a continuous, iterative process throughout the entire lifecycle of a high-risk AI System. All risks after mitigation, called residual risks, need to be judged acceptable (Soler Garrido et al., 2023).
Important in the context of non-discrimination law is Article 9(2)(a) AIA, which requires the identification and analysis of known and reasonably foreseeable risks that the high-risk AI system can pose to health, safety, or Fundamental Rights. As explained in Section 2.2, Fundamental Rights also include the right of non-discrimination. Thus, indirectly, Article 9 AIA already requires the identification and analysis of non-discrimination issues. Zuiderveen Borgesius et al. (2024) pointed out that this Article mandates the identification and mitigation of discrimination in AI systems. The initial draft of the AI Act did not include the reference to Fundamental Rights, and it was added by the EU Council in 2022 111111see Council of the European Union, “Proposal for a Regulation of the European Parliament and the Council laying down harmonized rules on artificial intelligence (Artificial Intelligence Act) and amending certain Union legislative acts – General approach” Doc. 14954/22, 25 November 2022. (Schuett, 2024), highlighting the need to ensure non-discriminatory regulations.
We identified three main problematic issues with the non-discrimination regulation in Article 9 AIA. First, only risks that are known and reasonably foreseeable need to be identified and analyzed. However, this is particularly problematic for Large Multimodal Models (LMMs) in high-risk contexts (Hacker et al., 2024a). What constitutes a reasonably foreseeable risk in an era where large language models are still being developed and not fully understood? Although alignment techniques have developed (Ji et al., 2023), discrimination still persists in LMMs (Gupta et al., 2023). When are these risks foreseeable? The AI Act leaves too much room for unclear speculation. From a technical perspective, benchmarks to judge the risk associated with LLMs have been proposed (Guldimann et al., 2024); however, automated bias testing is far from trivial, and we currently lack adequate ML techniques. Recital 65 AIA clarifies that the most appropriate risk-management measures should be used. There is a danger that even the most appropriate measures are not sufficient to identify foreseeable risks.
Second, the link to Fundamental Rights leaves many potential non-discrimination factors unaddressed (Kusche, 2024). According to Article 21(1) CFR, discrimination based on sex, race, color, ethnic or social origin, genetic features, language, religion or belief, political or any other opinion, membership of a national minority, property, birth, disability, age, or sexual orientation shall be prohibited. However, discrimination based on other factors is not prohibited. This creates unnecessary opportunities for discrimination. Furthermore, as explained above, the Fundamental Rights need to be balanced; the right of non-discrimination is only one among others. It is therefore unclear how effective this protection is (Kusche, 2024).
This issue is not unique to Article 9(2) AIA but is present in all Articles of the AI Act referring to Fundamental Rights. Health and safety are “bread and butter of product safety law” (Almada and Petit, 2023). The added link to Fundamental Rights adds a different dimension compared to product safety requirements. This mixture of aims is problematic since product safety law and Fundamental Rights law have different target regimes: technical aspects vs. individual rights.
Third, the relation between known and reasonably foreseeable risks and direct and indirect discrimination in traditional EU non-discrimination law is not well explored. Even if we assume that the risks through Fundamental Rights are indirectly related to non-discrimination requirements, it is unclear what known and reasonably foreseeable implies. Referring back to Section 2.2, traditional discrimination law differentiates between direct and indirect discrimination. Is (only) direct discrimination foreseeable? While indirect discrimination through algorithms seems more plausible, is it foreseeable? The exact relation is undefined.
Article 10(2)(f) AIA: Main input non-discrimination regulation of high-risk system. The core of the non-discrimination regulation for high-risk systems is Article 10 AIA. Article 10(2)(f) AIA requires that training, validation, and testing datasets shall be subject to an “examination in view of possible biases that are likely to affect the health and safety of persons, have a negative impact on fundamental rights or lead to discrimination prohibited under Union law, especially where data outputs influence inputs for future operations”.
Bias is a vague term. The first observation we make is that the input data of an AI system should be subjected to a bias analysis. Most importantly, the term bias is neither defined in the AI Act nor is there a common understanding of what bias means (see also Section 2.1) (van Bekkum, 2024). It seems that the regulators had a more technical definition of bias in mind, focusing on the diversity of training data in different dimensions compared to social, ethical, or structural biases (Hacker et al., 2024b). This makes it very hard to determine the regulatory content.
Do the negative effects need to be certain or likely? It is an open question to which condition the wording “that are likely” refers to. If the wording only refers to the first condition, then the negative impact on Fundamental Rights has to be certain. If the wording “that are likely” refers to the Fundamental Rights or to discrimination prohibited under Union law, then these conditions also include a probability statement. When comparing the wording of Article 10(2)(f) AIA to Article 9(2)(a) AIA, we note that Article 9(2)(a) targets the risk of Fundamental Rights impact. The risk is balancing between the likelihood of an occurrence and the severity of harm (Article 3(2) AIA). It therefore seems more consistent to assume that the term “likely” in Article 10(2)(f) AIA refers to all conditions. This view is also supported by the translation of the AI Act in different languages such as German, where it is more evident that the wording “that are likely” applies to all conditions.
Article 10(2)(f) AIA contrasts with Article 9(2)(a) AIA. In contrast to Article 10 AIA, the bias must have a likely negative impact on Fundamental Rights or lead to discrimination, while Article 9 uses the word “risk”. Again, risk is a combination of likelihood and severity. Compared to the risk in Article 9(2) AIA, in Article 10(2)(f) AIA, the severity of the harm is not taken into account, only the likelihood. This difference between Article 9(2) AIA and 10(2)(f) AIA creates unnecessary regulatory uncertainty. It remains unclear why the legislator opted for the difference in applicability between Article 9 AIA and Article 10 AIA.
Through the wording of Article 10(2)(f) AIA, the regulation directly links to traditional non-discrimination law. This condition covers many different aspects of discriminatory effects. In general, this achieves a strong protection objective. The relationship between direct and indirect discrimination remains unclear. Is indirect discrimination likely? Furthermore, intersectional aspects of discrimination are only vaguely covered by traditional discrimination law (see Sec. 2.2). Are they likely as well? When artificial intelligence systems use many different features as input, intersectional discrimination is likely (Wang et al., 2022; Roy et al., 2023).
Article 10(2)(f) AIA only targets the input data for machine learning. The examination is only mandatory for training, validation, and testing data. In other words, it only covers the input data of an AI System. Although the second half of the sentence reads “especially where data outputs influence inputs for future operations”, output data itself is not the main focus.
Furthermore, the wording of Article 10(2)(f) AIA implies an ex-post view: the developer has to know whether a bias will likely lead to discrimination. However, especially for larger models, it is technically very challenging and impossible to predict the individual output of AI models (Hacker et al., 2024b; Black et al., 2022).
Article 10(2)(g) AIA: Appropriate bias detection, prevention, and mitigation on input data must include factors beyond technical means. Article 10(2)(f) AIA is completed by Article 10(2)(g) AIA. Article 10(2)(g) AIA demands that “appropriate measures to detect, prevent and mitigate possible biases identified according to point (f)” are considered. Thus, Article 10 AIA not only mandates the examination of biases that lead to discrimination but also the mitigation of these biases. The bias tests must be documented and disclosed according to Article 11 AIA, along with Annexes IV and IXa AIA (Zuiderveen Borgesius et al., 2024). It was argued that with this focus on the input side, the AI Act seeks to remedy the root cause of biases (Zuiderveen Borgesius et al., 2024).
Input bias data detection, prevention, and mitigation methods remain largely unexplored. These methods are not clearly defined in the AI Act (Wachter, 2024; Deck et al., 2024b). Article 10(2)(g) AIA appears to be inspired by the technical literature on fairness metrics for algorithmic outputs. However, compared to output fairness metrics, technical bias detection methods for input data are understudied. Furthermore, as previously noted, the causes of biases in ML are multi-causal (Gebru, 2019; Gebru et al., 2021; Hanna et al., 2020; Mehrabi et al., 2021; Bender et al., 2021). Even when data are collected representatively, biases can occur (see (Mehrabi et al., 2021) for a comprehensive list). Two examples illustrate this argument:
Social bias can occur regardless of sample representativeness (Mitchell et al., 2021). For instance, in loan application settings, if certain groups (such as women) systematically receive lower wages, the data will contain inherent bias even with perfect representation.
Bias can emerge when one subgroup is smaller than others (Barocas et al., 2023). Intuitively, an algorithm’s performance decreases when trained on fewer data points, resulting in poorer performance for smaller subgroups. Since smaller subgroups often correspond to minorities, algorithms may perform worse for these populations even with representative data.
The data governance strategies in Article 10(2)(g) AIA focus too heavily on technical aspects of input data biases. While systematic fairness approaches have been developed for output bias detection, prevention, and mitigation (Han et al., 2023; Krasanakis and Papadopoulos, 2024; Pagano et al., 2023), we lack similarly comprehensive approaches for input data analysis. This creates a significant gap between the legal requirements in Article 10(2)(g) AIA and their technical and ethical feasibility.
Article 10(2) AIA: Additional obligations without specific fairness regulations In reviewing of Article 10(2) AIA, it is noticeable that the other sections (a-e) of the Article also indirectly link to algorithmic fairness considerations. For example, the representativeness of training data is explicitly mentioned in Article 10(2)(d). When interpreting Article 10(2)(f) AIA or Article 10(2)(g) AIA, however, the additional requirements in Article 10(2) AIA do not result in any new insights. Since these articles have no direct requirements regarding non-discrimination, we will not analyze them in further detail.
Article 10(5) AIA: A legal basis for the processing of personal data in non-discrimination contexts. Finally, Article 10(5) AIA of the AI Act allows the processing of Article 9 GDPR data (see also Recital 70 AIA). Article 9 GDPR protects special categories of personal data, such as genetic, biometric, or health data (Article 9(1) GDPR). There is tension (Deck et al., 2024a) between the need for debiasing AI algorithms and data protection law, which Article 10(5) tries to solve. In order to effectively mitigate biases in AI systems, the processing of personal data (for example, to compute fairness metrics) is important (van Bekkum, 2024). Notable here is that this exception only applies in the high-risk regime. Thus, for non-high-risk systems, this exception is not applicable. However, it may also be a legal basis for the processing of personal data outside of high-risk systems (Hacker, 2021).
Article 15 (4) AIA: Output bias regulation in case of feedback loops. In contrast to Article 10(2)(f) AIA, Article 15(4) AIA mandates that “High-risk AI systems that continue to learn after being placed on the market or put into service shall be developed in such a way as to eliminate or reduce as far as possible the risk of possibly biased outputs influencing input for future operations (feedback loops) and as to ensure that any such feedback loops are duly addressed with appropriate mitigation measures.” In this case, the output of a system is regulated, in contrast to the input of an AI system in Article 10 AIA. The reasoning behind this is that AI systems should not become echo chambers for biases (Zuiderveen Borgesius et al., 2024). According to Recital 67 AIA, this is a special concern when examining historical biases. Most important to note here is the fact, that this output regulation only applies if the systems continue learning after being placed on the market or put into service.
Additionally, it remains unclear if the word “bias” here refers to the same bias as in Article 10 AIA. Recital 67 AIA, which relates to Article 10 AIA but also mentions feedback loops, indicates that the same biases are addressed.
Biased output is usually detected and mitigated with fairness metrics. As shown above, the interpretation of fairness metrics is challenging. An example of this difficulty is presented in Figure 2 in the appendix. For some AI systems, there exists a trade-off between fairness and other measures such as accuracy of an AI system. Figure 2 shows a typical algorithmic fairness vs. accuracy trade-off for an AI system used to predict the income status, data are from(Mittelstadt et al., 2023). Although this trade-off can be reduced with approaches like “Levelling Up” (Mittelstadt et al., 2023), the trade-off is still present. Therefore, it becomes very challenging to identify how to “reduce as far as possible the risk of possibly biased outputs” in Article 15(4) AIA. This requirement is vague.
Furthermore, especially for language (generative) models, biases in the input data and biases in the output data might be different. Furthermore, the mention of bias in Article 15, titled “Accuracy, robustness, and cybersecurity”, is rather odd since the latter two fields are traditionally not mainly concerned with discrimination (Nolte et al., 2024). Here, it remains an open question whether the robustness of an AI system and fairness represent a trade-off (Lee et al., 2021).
Other articles regarding high-risk systems regulations. While the previously discussed articles provide at least some intuition about addressed concerns, we want to categorize the remaining articles of Table 1. Most of these paragraphs only concern Fundamental Rights impacts which show limited impact for non-discrimination regulation (see above). Thus, the discussion here is rather short. Article 13(3)(b)(iii) AIA mandates that information on the foreseeable misuse and Fundamental Rights impact is provided. Article 14(2) AIA mandates that human oversight is needed to minimize risks to Fundamental Rights. Article 17(1)(i) AIA and Article 26(5) AIA are related to reporting and information of serious incidents. Serious incidents are linked to discrimination through Article 3(49)(c) AIA: “’serious incident’ means an incident or malfunctioning of an AI system that directly or indirectly leads to any of the following: […] the infringement of obligations under Union law intended to protect Fundamental Rights”. The link to discrimination is even more indirect since it does not directly concern Fundamental Rights but instead Union law intended to protect Fundamental Rights. The relation to discrimination law becomes increasingly relaxed. Finally, Article 27(1) AIA requires a Fundamental Rights impact assessment when the high-risk system is used in special systems within bodies governed by public law. Yet again, significant gaps exist between theoretical and methodological elaboration of these impact assessments (Mantelero, 2024).
Summary: High-risk systems show vague non-discrimination regulations. Regulations regarding high-risk systems are vague. In contrast to the ambitions of the AI Act, the concrete non-discrimination regulations lack specificity. For the ML community it is challenging to understand the implications of the AI Act. Only Article 10 AIA and Article 15 AIA explicitly mention non-discrimination and biases, compared to the vague statement of Fundamental Rights protection in Article 9 AIA. However, Article 10 AIA only considers the input phase, while Article 15 AIA only addresses the output phase for feedback loops. It would be beneficial if Article 15 AIA, Article 9 AIA, and Article 10 AIA regulated the same types of direct, indirect, and intersectional discrimination on both the input and output sides.
3.4. General Purpose AI non-discrimination regulations
GPAI models are an important model category. Most (Multimodal) Large Language Models fall under this category. While most of our paper focuses on high-risk regulation in the fairness domain, we briefly discuss these important GPAI models as well. Besides the high-risk systems, GPAI systems are also heavily regulated (Novelli et al., 2024).
Introduction to GPAI models. Besides high-risk systems, the AI Act regulates so-called General Purpose AI (GPAI) systems in Article 51-56 AIA  [^12] Please note that the AI Act also regulates, as a third category, certain AI systems in Article 50 AIA.. According to Article 3(6) AIA, a GPAI system is “an AI system which is based on a general-purpose AI model and which has the capability to serve a variety of purposes, both for direct use as well as for integration in other AI systems”.
GPAI models are thus models with a wide variety of capabilities that are not immediately foreseeable after training and rely on model size (Wei et al., 2022). From a technical point of view, the definition includes Large Language Models as well as Large Multimodal Models such as OpenAI’s GPT-Model family, Anthropic’s Claude models, and Google Gemini (Triguero et al., 2024). If an AI System has GPAI capabilities as well as high-risk properties, it must adhere to both GPAI and high-risk regulations (Nolte et al., 2024).
Non-discrimination regulation in GPAI models with systemic risks. GPAI models and GPAI models with systemic risk must be differentiated. The AI Act also defines GPAI models with systemic risk in Article 51(1) AIA. A GPAI model with systemic risk is either a model that has high-impact capabilities evaluated on the basis of appropriate technical tools and methodologies, including indicators and benchmarks (Article 51(1)(a) AIA), or any GPAI model whose training compute exceeds $10^{25}$ floating point operations (flops). Especially, the last condition is a topic of debate (Heim and Koessler, 2024; Novelli et al., 2024; Gupta et al., 2024). Many researchers in ML try to build models with ChatGPT-like abilities but with less compute power, for example, through quantization (Xiao et al., 2023; Gholami et al., 2022; Ma et al., 2024). Thus, the definition of GPAI model with systemic risk based on training compute is questionable. The systemic risk itself is defined in Article 3(65) AIA. A systemic risk “means a risk that is specific to the high-impact capabilities of general-purpose AI models, having a significant impact on the Union market due to their reach, or due to actual or reasonably foreseeable negative effects on public health, safety, public security, fundamental rights, or the society as a whole, that can be propagated at scale across the value chain”. The risk must be specific to the high-impact capabilities of the GPAI model. All risks that are also present in low- to high-risk systems are not covered as systemic risk. Furthermore, the impact on Fundamental Rights — and thus on non-discrimination regulation — needs to be actual or reasonably foreseeable. In common with high-risk system regulations, not all risks need to be considered but only the actual or foreseeable risks. Adding the word “reasonable” before “foreseeable” softens the non-discrimination regulation. The basic regulations for normal GPAI models are given in Article 53 AIA, while the specific regulations for GPAI models with systemic risks are stated in Article 55 AIA. However, both articles do not include any specific regulations but rather quite vague documentation, evaluation, and analysis obligations for the providers. Although there are technical measures to assess the risks of GPAI models (Paskov et al., 2024; Barrett et al., 2023; Guldimann et al., 2024), it is currently rather unclear how to deal with the GPAI regulations. The legislator has already anticipated this issue. Therefore, Article 56 AIA calls for the EU AI Office to draw up a Code of Practice together with providers and other stakeholders. Currently, this Code of Practice is being developed in four directions: Transparency and copyright-related rules, Risk identification and assessment, Technical risk mitigation, and Internal risk management and governance of General-purpose AI providers (Cod, 2024). It currently remains open whether this will result in more concrete regulations that enable technical and legal compliance.
Summary: GPAI models are only vaguely regulated. In summary, GPAI models are currently only vaguely regulated. There exists significant regulatory uncertainty, especially regarding how the regulations of the AI Act can be implemented in practice. Perhaps the Code of Practice being developed by the AI Office and the European Commission can provide guidance in this area.
4. Summary & Outlook: How to proceed?
As Mayson (2018) already pointed out: Bias in, bias out. The AI act seek to complement existing non-discrimination EU law with it processed based regulation (Zuiderveen Borgesius et al., 2024). However, even on this broader level, the AI Act fails to account for major issues.
First, the AI Act heavily focuses on discrimination because of data. The non-discrimination regulations analyzed in the above section mostly focus on bias in (training) data and thus exclude other sources of bias, such as design choices in the algorithms. Furthermore, it is missing specificity on detection and mitigation methods. That makes it hard for practitioners to follow the law.
Second, the AI Act does not include any regulations for none to limited-risk systems. Many AI systems are not regulated at all within the AI Act. It is worth noting that the EU estimated that high-risk systems comprise roughly 5%-15% of all AI systems in the EU. Although the remaining 85%-95% of all systems include a few GPAI systems, most AI systems remain entirely uncovered. Non-model specific regulations, such as the obligation to foster AI literacy in Article 4 AIA, are insufficient to ensure non-discriminatory practices. From our perspective, the legislator is missing an opportunity to regulate AI systems on a broader scope. This concerns not only the obligations but also specific rights. For example, the use of personal data for bias detection could also be beneficial for other than high-risk systems.
Many of the rules focus on (internal) compliance (Zuiderveen Borgesius et al., 2024) and checklist rather than on output-based safeguards for users affected by the AI system This internal compliance is also linked to the self-assessment approach. It is questionable whether the self-assessment is suitable for protecting individuals and their Fundamental Rights as well as non-discrimination regulation (Bosoer et al., 2023).
Future: A call for more interdisciplinary work. While the AI Act is casuistic and incomplete at the same time (Ződi, 2024), we believe that the AI Act would benefit from four refinements. (1.) If the legislators decide to target non-discrimination regulation, non-discrimination should be explicitly mentioned. In contrast to vague terms such as bias, Fundamental Rights, or fairness, non-discrimination is a well-established legal concept within other European legislative frameworks. (2.) The non-discrimination regulation should cover input as well as output parts of the algorithms in one single article, also along the full AI life cycle (Deck et al., 2024b). Algorithmic output in terms of non-discrimination is only regulated if systems continuously learn after being placed on the market. Otherwise, outputs in terms of discriminatory results of AI systems are not regulated at all. Technical and non-technical means for bias detection, prevention, and mitigation for input and output data should be developed. (3.) Clarification is needed on appropriate measures for non-discrimination. While a reverse burden-of-proof approach is possible here (Wachter, 2024), similar processes as the Code of Practice for GPAI models might be possible such that technical and legal experts as well as all stakeholders work together on refined non-discrimination measures. (4.) If GPAI models are regulated in the current form, a closer link between GPAI models and downstream applications needs to be established (Hacker et al., 2024b). Responsibilities for non-discrimination law compliance must be specified. It is especially important to note that this work is best done in joint community efforts of legal researchers and ML researchers. A bridge between these two fields is highly desirable since regulation should be guided by technical appropriateness and feasibility.
In the introduction of our work, we raised how fair results of AI algorithms can be ensured. We showed that the AI Act tried to solve this non-trivial task with a normative non-discrimination approach. This paper aimed to address legal researchers with an interest in Machine Learning as well as computer scientists and related domains with an interest in legal regulations. We intend to bridge the gap between regulatory and technical limitations in the machine learning domain. From our perspective, both disciplines can benefit from tighter interactions.
5. Acknowledgments
We would like to thank Miriam Rateike for her detailed comments on an earlier version of the manuscript. We are grateful for comments from Christoph Kern and the Social Data Science and Statistical Learning group during an institute seminar, which clearly improved this paper. Furthermore, we thank Alina Wenick, Tharushi Abeynayaka and Mara Seyfert for comments on the manuscript and proofreading the paper.
Kristof Meding was supported by the Carl Zeiss Foundation through the CZS Center for AI and Law. Additionally, Kristof Meding is a member of the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645.