### v3

# It’s complicated. The relationship of algorithmic fairness and non-discrimination regulations for high-risk systems in the EU AI Act

Abstract
What constitutes a fair decision? This question is not only difficult for humans but becomes more challenging when Artificial Intelligence (AI) models are used. In light of discriminatory algorithmic behaviors, the EU has recently passed the AI Act, which mandates specific rules for high-risk systems, incorporating both traditional legal non-discrimination regulations and machine learning based algorithmic fairness concepts. This paper aims to bridge these two different concepts in the AI Act through: First, a necessary high-level introduction of both concepts targeting legal and computer science-oriented scholars, and second, an in-depth analysis of the AI Act’s relationship between legal non-discrimination regulations and algorithmic fairness. Our analysis reveals three key findings: (1.) Most non-discrimination regulations target only high-risk AI systems. (2.) The regulation of high-risk systems encompasses both data input requirements and output monitoring, though these regulations are partly inconsistent and raise questions of computational feasibility. (3.) Finally, we consider the possible (future) interaction of classical EU non-discrimination law and the AI Act regulations. We recommend developing more specific auditing and testing methodologies for AI systems. This paper aims to serve as a foundation for future interdisciplinary collaboration between legal scholars and computer science-oriented machine learning researchers studying discrimination in AI systems.
1 Introduction
How can we ensure fair algorithms in the context of AI systems? When regulating algorithms and, more specifically, Artificial Intelligence (AI) systems, this question becomes fundamental. While the question is also crucial in an analog world, it becomes increasingly pressing in the digital world. The European Union (EU) has become one of the forerunners of regulations in the digital age with regulatory frameworks, such as the Digital Services Act [^1] Regulation (EU) 2022/2065., the Digital Markets Act [^2] Regulation (EU) 2022/1925., or the Data Act [^3] Regulation (EU) 2023/2854.. In order to maintain fair data processing in the machine learning domain, the European Union (EU) recently adopted the Artificial Intelligence Act (AIA) [^4] Regulation (EU) 2024/1689.. The aim of this AI regulation is to maintain a level playing field for “ethical principles” (Nolte, Rateike, and Finck 2024) within and outside the European Union (see recital 27 AIA).
Fair algorithmic processing also matters from a computational perspective. Over the past decade, algorithmic fairness has become a well-established field within machine learning that focuses on defining, mitigating, and evaluating discriminatory behavior of Artificial Intelligence models (Pessach and Shmueli 2023). The importance becomes evident when looking at several discriminatory behaviors of algorithms in the past, ranging from hiring (Dastin 2022) to social welfare systems (Hadwick and Lan 2021). In recent years, large generative (multimodal) models, particularly Large Language Models (LLMs) with their high accessibility and wide range of applications have posed significant new challenges to the algorithmic fairness domain (Chu, Wang, and Zhang 2024; Kotek, Dockum, and Sun 2023).
The relationship between computer science and legislation within the AI Act is complex. While the interplay between algorithmic fairness and non-discrimination law is often unclear (Weerts et al. 2023), efficient AI legislation requires interdisciplinary collaboration connecting the legal domain and computer science. Although legal and technical concerns regarding algorithmic fairness and non-discrimination law have existed for a long time (Barocas and Selbst 2016; Hacker 2018; Wachter, Mittelstadt, and Russell 2021), recent developments have accelerated the pace. An increasing number of companies and laypersons are using these technologies, and with increased usage, the likelihood of misuse also rises. Thus, the need for efficient and effective regulations to protect individuals is substantial.
Previous literature. Discrimination through algorithms is not a novel phenomenon and was analyzed in a pre-AI Act era. Already in the 1980s, algorithmic discrimination was observed in admissions settings (Connors et al. 1981; Williams, Werth, and Wolff 1981), and gender inequalities in educational software were discussed (Huff and Cooper 1987). In 1996, researchers pointed out the levels at which such technical constraints and social institutions’ biases can occur (Friedman and Nissenbaum 1996). More recently, the interaction between AI and automated decision-making systems with traditional European and US non-discrimination law has been studied extensively. Barocas and Selbst (2016) presented a taxonomy of different sources of discrimination and their impact on humans. The relationship among various European non-discrimination laws has also been studied (Wachter, Mittelstadt, and Russell 2021, 2020; Hacker 2018; Weerts et al. 2023; Xenidis and Senden 2020; Lewis et al. 2025). For example, the seminal study by Wachter, Mittelstadt, and Russell (2021) presented a fairness metric that connects algorithmic fairness to the jurisprudence of the European Court of Justice in relation to the General Data Protection Regulation. Similarly, the connection between non-discrimination law and the GDPR has been analyzed to “unlock the algorithmic black box” (Hacker 2018). Naturally, since these papers were published before the AI Act was passed, they do not contain references to it.
Recently, several papers have already addressed the AI Act. Short studies such as Deck et al. (2024a); Ruohonen (2024) have analyzed the relationship between fairness and the AI Act. Additionally, sections within papers of broader scope beyond pure non-discrimination analyses have mentioned aspects of the relationship between algorithmic fairness and the AI Act (Wachter 2024; Hacker et al. 2024; Novelli et al. 2024). The AI Act and its relation to gender equality and non-discrimination law have also been discussed (Lütz 2024). Among these, the work by Bosoer, Cantero Gamito, and Rubio-Marin (2023) is most similar to ours in terms of its focus on the AI Act; however, it investigates non-discrimination regulations in the draft of the AI Act without a strong focus on the interaction with computer science. Most similar in terms of the interaction between computer science and legal research is the paper of Weerts et al. (2023). In contrast to our work, they did not focus on the AI Act.
Our contributions. This paper aims to foster and extend an interdisciplinary view on explicit and implicit fairness regulations in the AI Act regarding algorithmic fairness in the ML domain. We aim to introduce the field of algorithmic fairness to legally oriented researchers and discuss the fragmented relationship between the two fields. These sections are necessary for a broader audience to follow our reasoning.
For the in-depth analysis of the EU AI Act’s non-discrimination regulations in Section 3, our contributions are as follows
  1. 1.
    We present the history, scope, and intentions of non-discrimination regulations in the AI Act
  2. 2.
    We find that high-risk AI systems (Article 6 AIA) are most thoroughly regulated regarding non-discrimination. We discuss the relation between high-risk systems’ regulations and algorithmic fairness. We find that the AI Act mainly regulates input data compared to the output data of AI systems.
  3. 3.
    We discuss challenges for classical EU non-discrimination at the intersection of the AI Act, since classical EU non-discrimination law will still be important in AI contexts.
In the last section, we summarize our findings and give recommendations for future research. Additionally, based on our results, we suggest approaches to overcome some challenges and propose possible actions to address them.
2 A primer on non-discrimination from an AI and a legal perspective
Interdisciplinary research on algorithmic fairness is challenging, both from a computer science and a non-discrimination law perspective. Understanding legal reasoning can be challenging without prior legal knowledge, just as understanding algorithmic methods is difficult without a computational background.
However, the complexity in interdisciplinary work extends beyond pure knowledge. The challenges begin with terminology: For instance, in computer science, “fairness” is a term that can refer to different desiderata that aim to prevent socially or morally undesirable behavior or outcomes of algorithms. Although fairness has been discussed as a principle in economic law contexts (Scheuerer 2023), it is not a specific legal term. The arguably closest legal counterpart term is “non-discrimination”, which focuses on preventing unfair treatment based on characteristics such as race, gender, or other attributes on legal grounds. The interaction between these key concepts from law and computer science is demanding.
Thus, we highlight the differences and common grounds of algorithmic fairness and non-discrimination regulations in the following sections. After presenting an introduction to algorithmic fairness — for legal researchers — and an introduction to EU non-discrimination law — for computer science oriented researchers —, we attempt to bridge the gap between these two areas.
Algorithmic fairness: A gentle introduction
Algorithmic fairness has emerged as an established computer science and AI-related field over the past years. Algorithmic fairness focuses on uncovering and rectifying disadvantageous treatment of individuals or groups in machine learning models (Mitchell et al. 2021). This treatment must be considered within appropriate social, theoretical, and legal contexts (Pessach and Shmueli 2023) and includes evaluation and auditing methods to test for unfairness.
Origins of unfairness in algorithms. Unfair algorithmic behavior can have different sources. In order to discuss the origins, another — not well-defined — a term from computer science is often used: bias. There exist many different bias definitions. Since the EU AI Act does not define it itself, we use a definition from the EU Commission from 2021. According to this, a bias can be defined as “[…] bias describes systematic and repeatable errors in a computer system that create unfair outcomes, such as favouring one arbitrary group of users over others” (Commission 2021). Biases are primarily discussed within the data used for AI models and can be sorted into different categories. Most researchers emphasize the different aspects of data inequalities when discussing algorithmic fairness and sort data biases into different categories, such as the over-representation of specific groups in datasets (Zou and Schiebinger 2018). Hacker (2018) categorizes the biases in data into two groups: biased training data versus unequal ground truth; other authors use more categories. Barocas and Selbst (2016) uses five categories to map biases, while Mehrabi et al. (2021) uses 21 different categories for biases. It is important to note that the data are only one — albeit important — source of biases. For example, design decisions in the algorithms or structural power asymmetries also lead to algorithmic unfairness (Mehrabi et al. 2021; Gebru et al. 2021; Sousa e Silva 2024).
Approaches to algorithmic fairness: Quantification. Based on the different understandings of unfairness and for the purpose of uncovering and rectifying unjustified treatment between groups in algorithms, so-called fairness metrics have been proposed (Corbett-Davies et al. 2023; Castelnovo et al. 2022; Verma and Rubin 2018). The core idea of fairness metrics is to quantify algorithmic outputs and make them comparable through numerical measurements. Most metrics try to assess whether an output of an ML system is unfair to individuals (Dwork et al. 2012) or groups of people (Barocas, Hardt, and Narayanan 2023; Binns 2020). They provide a quantitative measure intended to indicate whether an algorithm demonstrates unjustified unequal treatment. Fairness metrics also include moral norms (Deck et al. 2024a; Hellman 2020).
The emergence of the field of algorithmic fairness has challenged existing approaches to ML by highlighting the need to contextualize them, decide on, and provide a rationale for the optimization criteria chosen. First, it has been shown that an apparently objective number needs to be contextualized to the application under consideration (Wachter, Mittelstadt, and Russell 2021). Furthermore, some metrics are, from a mathematical viewpoint incompatible with each other [^5] They are also incompatible from a moral point of view (Heidari et al. 2019)., thus these fairness metrics cannot be fulfilled at the same time (Chouldechova 2017; Kleinberg, Mullainathan, and Raghavan 2016). This can lead to an effect of “d-hacking” or “fairness hacking” where users can choose their favorite metric to create the (technical and mathematical) impression that the algorithm is fair (Black, Gillis, and Hall 2024; Meding and Hagendorff 2024). Fairness metrics for mitigation can play a role at different stages of the development, testing and deployment of an algorithm. The pre-processing stage, the in-training stage, and the post-processing phases are differentiated (Barocas, Hardt, and Narayanan 2023; Binns 2018). In the pre-processing phase, the input data itself is altered to ensure fair processing (Caton and Haas 2024; Kamiran and Calders 2012). For the in-training phase, the optimization process during the training of an ML model is adjusted to include fairness as an optimization goal. Finally, in the post-processing phase, the output of an already trained model is adapted. The first and the last approaches make it possible to perform fairness analyses even if one only has black box access to the model (Caton and Haas 2024).
The era of LLMs raises new questions. LLMs have the advantage — or disadvantage, depending on the viewpoint — that their exact use case is most of the time not predefined, and they are trained on large amounts of various data. This diversity makes them later applicable to different contexts, from providing cookie recipes to coding exercises. However, this diversity also includes variations in unfairness. Algorithmic fairness in relation to LLMs is thus somewhat different from algorithmic fairness in classical AI (Doan et al. 2024; Chu, Wang, and Zhang 2024). Some researchers have applied classical fairness metrics in classification settings to LLMs (Chhikara et al. 2024). Additionally, it has been shown that LLM-specific issues arise due to the use of massive data and processing (Kotek, Dockum, and Sun 2023; Ferrara 2023; Navigli, Conia, and Ross 2023; Huang et al. 2024) at various steps of algorithmic development. Thus, it was argued that LLMs cannot yield fair outcomes at all (Anthis et al. 2024).
EU non-discrimination Law: A gentle introduction
We first describe fundamental rights as well as traditional EU non-discrimination law. Both influence the AI Act.
Fundamental rights in EU law. The Charter of Fundamental Rights of the European Union (CFR) is the main fundamental rights regulation within EU law [^6] The CFR should not be confused with European Convention on Human Rights (ECHR). The ECHR was not developed by the EU. Nevertheless, the EU needs to accede to the ECHR. The exact relationship of the ECHR to EU law is complex, see (Brittain 2015) for further discussion.. Through Article 6(1) of the Treaty on European Union, the Charter is part of the primary law of the EU.
The CFR includes non-discrimination law. For non-discrimination setting obligations, Article 21 CFR states: “Any discrimination based on any ground such as sex, race, color, ethnic or social origin, genetic features, language, religion or belief, political or any other opinion, membership of a national minority, property, birth, disability, age or sexual orientation shall be prohibited.” Thus, only discrimination by these listed attributes is targeted.
However, the right to non-discrimination is not absolute. It is important to note that the right to non-discrimination is only one right in the CFR, among others, such as the freedom of expression and information (Article 11(1) CFR). In order to apply the CFR, rights need to be balanced (Article 52(1) CFR). First, legislators must balance different CFR rights proportionally when creating new laws. Additionally, when laws reference the CFR, implementers must also follow this balancing obligation. A logic of proportionality of the different rights guides fundamental rights assessments (Almada and Petit 2023). Fundamental rights are not absolute, hierarchical, or quantifiable and need to be applied case by case (Sousa e Silva 2024). The provisions of the CFR target public parties (vertical applicability)(Fornasier 2015). Therefore, there are debates about how exactly the CFR applies to relationships between private parties (horizontal applicability) in general (Fornasier 2015; Cherednychenko 2007; Frantziou 2015; Prechal 2020) and in the AI Act context specifically (Lewis et al. 2025). In the context of the GDPR, scholars and the European Court of Justice (ECJ) noted that the CFR right can have a horizontal effect if the secondary law reflects a general principle of EU law (Ufert 2020). Since the AI Act also reflects general principles of EU law, we do not exclude some horizontal applicability of the CFR. When discussing fundamental rights in the AI Act, it is also important to keep the political and institutional landscape in mind (Palmiotto 2025).
Direct and indirect discrimination in EU Law. While the CFR is part of the primary law, it also influences the law of the European Union (Lewis et al. 2025). EU law has a long history of non-discrimination regulations and different laws include non-discrimination regulations [^7] See for an overview: https://commission.europa.eu/aid-development-cooperation-fundamental-rights/your-fundamental-rights-eu/know-your-rights/equality/non-discrimination˙en.. While also primary law include non-discrimnation regulations, one example is the prohibition of unequal payment based on gender (Article 157(1) TFEU), the focus of traditional non-discrimnation law in the EU is on secondary law (directives or regulations). For example, the Directive 2000/43/EC (Race Equality Directive) and Directive 2000/78/EC (Employment Equality Directive) include specific non-discrimination regulations.
In order to legally assess discrimination under the specific secondary EU Law, traditionally two different versions of discrimination are differentiated: direct and indirect discrimination (Zuiderveen Borgesius et al. 2024; Wachter, Mittelstadt, and Russell 2020). However, importantly, EU-based direct discrimination does not require any intentional wrongdoing (Weerts et al. 2023; Xenidis and Senden 2020; Adams-Prassl, Binns, and Kelly-Lyth 2023).
Direct discrimination is defined as situations in which “one person is treated less favourably than another is, has been or would be treated in a comparable situation” (Article 2(2)(a) Directive 2000/43/EC). This means that individuals are discriminated against by using a (listed) protected attribute.
Indirect discrimination is much more challenging to tackle. In indirect discrimination cases, seemingly neutral attributes are used, but they rely on a protected attribute. For indirect discrimination, it is important to note that it can be justified through legitimate aims and appropriate means (Zuiderveen Borgesius et al. 2024).
Also, a third category of discrimination needs to be taken into account [^8] Please note that other non-discrimination rules in sector-specific regulations, such as the EU Consumer Credit Directive, are beyond the scope of our paper., which challenges traditional non-discrimination law: intersectional discrimination. Intersectional discrimination concerns cases “originating in several inextricably linked vectors of disadvantage” (Xenidis 2023). This becomes especially important since (modern) AI models do not use single variables as input — such as our simplified loan example based on the single variable above — but instead use many different aspects as input. This can lead to effects where discrimination may only occur at the intersection of gender and age(Weerts et al. 2023). It is an open question if and how, at the current stage, the European Court of Justice recognizes intersectional discrimination as a form of discrimination (Weerts et al. 2023; Xenidis 2023; Atrey 2018).
3 EU AI Act’s non-discrimination regulations for high-risk systems.
Before we discuss the specific non-discrimination regulations for high-risk systems, we will briefly introduce the history, scope, and most important definitions of the AI Act. This short introduction is followed by an overview of the AI Act’s non-discrimination regulations for high-risk systems.
The EU AI Act: History, Scope, and Definitions
In this section, we first describe the origins of the EU AI Act and clarify key concepts such as risk, systems, and the difference between developers and deployers, which are necessary to understand our analysis of the AI Act.
Emergent technology can benefit from efficient regulation. As one of the first comprehensive AI regulations worldwide (Wodi 2024), the AI Act (AIA) introduces harmonized rules and obligations for the use and “placing on the market” of AI systems within the European Union (Article 1(2) AIA). It is primarily a product safety regulation (Almada and Petit 2023), aiming to establish a level playing field for AI technologies across the Union (Nolte, Rateike, and Finck 2024). Notably, the original drafts of the EU Commission did not include explicit provisions on individual rights (Hacker et al. 2024). At this stage, individual rights — understood as protecting the rights of individuals affected — were largely absent. In fact, Members of the European Parliament initially did not prioritize the regulation of algorithmic discrimination (Chiappetta 2023). However, over the course of the legislative process, different aspects of non-discrimination regulations were integrated into the final text.
The AI Act regulates AI systems and also includes provisions for General Purpose AI (GPAI) models. The relationship or distinction between a model and a system in the context of the AIA remains legally unresolved. Article 3(1) defines an “AI system” as “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments“, but the article does not provide a definition of “AI model”. Recital 97 clarifies that “although AI models are essential components of AI systems, they do not constitute AI systems on their own. AI models require the addition of further components, such as, for example, a user interface, to become AI systems.” This formulation does not conclusively settle the terminological relationship between models and systems, as discussed in more detail in Nolte, Rateike, and Finck (2024). Moreover, the regulation of “AI systems” contrasts with the (trustworthy) computer science literature, which typically studies machine learning models. Since the focus here is on the legal framework, we adopt the terminology used in the AI Act.
The core regulatory structure of the AI Act is built on a risk-based approach (Hacker, Cordes, and Rochon 2024). Risk is defined in Article 3(2) AIA as “the combination of the probability of the occurrence of harm and the severity of that harm.” Based on this definition, the AI Act categorizes AI systems into different risk levels, each associated with a specific set of regulatory requirements [^9] This classification has been subject to criticism, see e.g., Bosoer, Cantero Gamito, and Rubio-Marin (2023).:
  • Unacceptable risk (Art. 5 AIA): These AI systems are prohibited. Examples include social scoring by governments.
  • High-risk (Art. 6ff. AIA): These systems are subject to stringent obligations, such as requirements for robustness, accuracy, or non-discrimination. Examples include AI used in recruitment, credit scoring, or law enforcement.
  • Certain AI systems (Art. 50 AIA) with specific risk: These systems must meet transparency obligations, such as informing users they are interacting with an AI system. Examples include systems that produce deep-fakes.
  • All other systems: These systems are not subject to specific regulatory obligations under the AIA, except broad regulations like Art. 4 AIA (AI Literacy).
In addition, the AIA includes specific provisions for General Purpose AI (GPAI) models (Articles 51 ff. AIA). The use case is not predefined, which is specific to GPAI models. LLMs are an example of a classical GPAI model. The AI Act distinguishes between general-purpose models posing “systemic risk” and those that do not.
The primary addressees of the AI Act are providers (Article 3(3) AIA) and deployers (Article 3(4) AIA). Providers are the entities responsible for developing and placing an AI model on the market or putting an AI system into service. In contrast, deployers are those who use an AI system under their authority. Many of the obligations we discuss in the following sections primarily concern the providers of AI systems, while some apply specifically to deployers (e.g., Article 26 AIA).
Non-discrimination regulations within the AI Act.
Analyzing non-discrimination regulations starts with identifying regulations. In total, we scanned the AI Act for non-discrimination-related terms: discrimination, fundamental right, fairness, and bias. We noticed that within the definitions of Article 3 AIA, the terms serious incident (Article 3(49) AIA) and systemic risk (Article 3(65) AIA) refer to fundamental rights. Thus, these were included in our analysis as well.
While analyzing the regulations, we observed that terms appear in the majority of cases in the recitals, rather than in the articles of the AI Act. The recitals are only considered helpful when interpreting the articles (Klimas and Vaiciukaite 2008) but do not have a legally binding force. Therefore, our analysis does not focus on the recitals. A full table with all articles of the AI Act, including non-discrimination-related terms, can be found in the appendix in Table 1.
This analysis shows, first, that the majority of non-discrimination regulations in the EU AI Act concern the regulation of high-risk systems. GPAI models are only implicitly regulated by the systemic risk and serious incident terms. This will have an impact on our further analysis: we focus on high-risk systems compared to GPAI models.
Second, most non-discrimination regulations are only implicitly addressed by the reference to fundamental rights. Discrimination and bias as threats are explicitly addressed only for high-risk systems. Here too, we observe a relation to fundamental rights.
Third, the term “fairness” from algorithmic fairness is not used at all within the AI Act. On the one hand, this result is surprising. The absence of the fairness aspect reveals a gap between the wording used in the computer science and ML communities and that of the legal field. On the other hand, however, as the non-discrimination rules are based on established rules within EU legislative frameworks, it is not a surprising finding.
After analyzing which parts of the AI Act contain non-discrimination regulations, we will now analyze the details for high-risk systems in detail.
High-risk AI non-discrimination regulations.
What is a high-risk AI system? Article 6 AIA defines different aspects of high risks. In combination with Annex III, high-risk systems are, for example, intended to be used as safety components in the management and operation of critical digital infrastructure (Annex III(2) AIA) or systems intended to be used for the recruitment or selection of natural persons (Annex III(4)(a) AIA). The reasoning behind the specific regulations in Article 6 AIA is that systems which pose a specific risk to fundamental rights or safety need tighter regulation (Fraser and y Villarino 2024) than other less risky systems. Recital 48 states that the adverse impact caused by the AI system on fundamental rights is of particular relevance when classifying an AI system as a high-risk system. It is expected that 5%-15% [^10] For a critical assessment of these numbers see (Almada and Petit 2023) of all AI systems fall under the category of high-risk systems (Commission 2021).
We identified Articles 9, 10, and 15 AIA as the main non-discrimination provisions of high-risk systems. These will be discussed in detail in the following paragraphs [^11] Please note that we do not differentiate here between the obligations for providers and those obligations for the deployers of AI systems, since we focus on the interaction of computer science and law.. Furthermore, Article 11 (together with Annex III) and Article 13 AIA will be discussed due to their relationship to the aforementioned articles.
Article 9 AIA: Non-discrimination regulations within risk management systems Article 9 of the AI Act covers risk management systems of high-risk systems. According to Article 9(2)(a-d) AIA, risk management systems are designed to identify, evaluate, and mitigate risks of AI systems in a continuous, iterative process throughout the entire lifecycle of a high-risk AI system. All risks after mitigation, called residual risks, need to be judged acceptable (Soler Garrido et al. 2023).
Important in the context of non-discrimination law is Article 9(2)(a) AIA, which requires the identification and analysis of known and reasonably foreseeable risks that the high-risk AI system can pose to health, safety, or fundamental rights. Through the link to fundamental rights, indirectly, Article 9 AIA already requires the identification and analysis of non-discrimination issues (Zuiderveen Borgesius et al. 2024). The initial draft of the AI Act did not include the reference to fundamental rights, and it was added by the EU Council in 2022  [^12] See Council of the European Union, “Proposal for a Regulation of the European Parliament and the Council laying down harmonized rules on artificial intelligence (Artificial Intelligence Act) and amending certain Union legislative acts – General approach” Doc. 14954/22, 25 November 2022..
We identified three challenges from an algorithmic fairness perspective with the non-discrimination regulation in Article 9 AIA. First, only risks that are known and reasonably foreseeable need to be identified and analyzed. However, this is particularly problematic for Large Multimodal Models (LMMs) in high-risk contexts (Hacker, Cordes, and Rochon 2024). What constitutes a reasonably foreseeable risk in an era where large language models are still being developed and not fully understood? Although alignment techniques have been developed (Ji et al. 2023), discrimination still persists in LMMs (Gupta et al. 2023). When are these risks foreseeable? The AI Act leaves much room for speculation. From a technical perspective, benchmarks to judge the risk associated with LLMs have been proposed (Guldimann et al. 2024); however, automated bias testing is far from trivial, and we currently lack adequate ML techniques. Recital 65 AIA clarifies that the most appropriate risk-management measures should be used. There is a danger that even the most appropriate measures are not sufficient to identify foreseeable risks.
Second, the link to fundamental rights leaves potential non-discrimination factors unaddressed (Kusche 2024). According to Article 21(1) CFR, discrimination based on sex, race, color, ethnic or social origin, genetic features, language, religion or belief, political or any other opinion, membership of a national minority, property, birth, disability, age, or sexual orientation shall be prohibited. However, discrimination based on other factors is not prohibited by Article 21(1) CFR but may be prohibited by other European or national laws. This creates unnecessary opportunities for discrimination. Furthermore, as explained above, the fundamental rights need to be balanced; the right of non-discrimination is only one among others. It is therefore unclear how effective this protection is (Kusche 2024).
This issue is not unique to Article 9(2) AIA but is present in all articles of the AI Act referring to fundamental rights. Health and safety are “bread and butter of product safety law” (Almada and Petit 2023). The added link to fundamental rights adds a different dimension compared to product safety requirements. This mixture of aims is problematic since product safety law and fundamental rights law have different target regimes: technical aspects vs. individual rights.
Third, the relation between known and reasonably foreseeable risks and direct and indirect discrimination in traditional EU non-discrimination law is not well explored (Zuiderveen Borgesius et al. 2024; Wachter, Mittelstadt, and Russell 2020). Even if we assume that the risks through fundamental rights are indirectly related to non-discrimination requirements, it is unclear what known and reasonably foreseeable implies. Is (only) direct discrimination foreseeable? While indirect discrimination through algorithms seems more plausible, is it foreseeable? The exact relation is undefined.
Article 10(2)(f) AIA: Main input non-discrimination regulation of high-risk system. The core of the non-discrimination regulation for high-risk systems is Article 10(2)(f) AIA. Article 10(2)(f) AIA requires that training, validation, and testing datasets shall be subject to an “examination in view of possible biases that are likely to affect the health and safety of persons, have a negative impact on fundamental rights or lead to discrimination prohibited under Union law, especially where data outputs influence inputs for future operations”.
The first observation we make is that the input data of an AI system should be subjected to a bias analysis. Most importantly, the term bias is neither defined in the AI Act nor is there a common understanding of what bias means (van Bekkum 2024). It seems that the regulators had a more technical definition of bias in mind, focusing on the diversity of training data in different dimensions compared to social, ethical, or structural biases (Hacker et al. 2024). This could imply difficulties in determining the regulatory content.
Do the negative effects need to be certain or likely? It is an open question to which condition the wording “that are likely” refers to. If the wording only refers to the first condition, then the negative impact on fundamental rights has to be certain. If the wording “that are likely” refers to the fundamental rights or to discrimination prohibited under Union law, then these conditions also include a probability statement. When comparing the wording of Article 10(2)(f) AIA to Article 9(2)(a) AIA, we note that Article 9(2)(a) targets the risk of fundamental rights impact. The “risk” is a balance between the likelihood of an occurrence and the severity of harm (Article 3(2) AIA). It therefore seems more consistent to assume that the term “likely” in Article 10(2)(f) AIA refers to all conditions. This view is also supported by the translation of the AI Act in different languages such as German, where it is more evident that the wording “that are likely” applies to all conditions.
Article 10(2)(f) AIA contrasts with Article 9(2)(a) AIA. In contrast to Article 10 AIA, the bias must have a likely negative impact on fundamental rights or lead to discrimination, while Article 9 uses the word “risk”. Again, risk is a combination of likelihood and severity. Compared to the risk in Article 9(2) AIA, in Article 10(2)(f) AIA, the severity of the harm is not taken into account, only the likelihood. This difference between Article 9(2) AIA and 10(2)(f) AIA creates regulatory uncertainty. It remains open why the legislator opted for the difference in applicability between Article 9 AIA and Article 10 AIA.
Through the wording of Article 10(2)(f) AIA, the regulation directly links to traditional non-discrimination law. Many different aspects of discriminatory effects are covered. On the one hand, the link achieves a strong protection objective for the input side of the AI systems. However, even without this link, Union law, which prohibits discrimination, could be applicable (see also next section). Thus, the question remains: what new legislative content is added? On the other hand, all existing problems within traditional non-discrimination law remain. For example, the relationship between direct and indirect discrimination remains unclear. Is indirect discrimination likely? Furthermore, intersectional aspects of discrimination are only vaguely covered by traditional discrimination law. Are they likely as well? When artificial intelligence systems use many different features as input, intersectional discrimination is likely (Wang, Ramaswamy, and Russakovsky 2022; Roy, Horstmann, and Ntoutsi 2023).
Article 10(2)(f) AIA only targets the input data for machine learning. The examination is only mandatory for training, validation, and testing data. In other words, it only covers the input data of an AI system. Although the second half of the sentence reads “especially where data outputs influence inputs for future operations”, output data itself is not the main focus.
Furthermore, the wording of Article 10(2)(f) AIA implies an ex-post view: the developer has to know whether a bias will likely lead to discrimination. However, especially for larger models, it is technically very challenging and impossible to predict the individual output of AI models (Hacker et al. 2024; Black, Raghavan, and Barocas 2022).
Article 10(2)(g) AIA: Appropriate bias detection, prevention, and mitigation on input data must include factors beyond technical means. Article 10(2)(f) AIA is complemented by Article 10(2)(g) AIA. Article 10(2)(g) AIA demands that “appropriate measures to detect, prevent and mitigate possible biases identified according to point (f)” are considered. Thus, Article 10 AIA not only mandates the examination of biases that lead to discrimination but also the mitigation of these biases. The bias tests must be documented and disclosed according to Article 11 AIA, along with Annexes IV and IXa AIA (Zuiderveen Borgesius et al. 2024). It has been argued that with this focus on the input side, the AI Act seeks to remedy the root cause of biases (Zuiderveen Borgesius et al. 2024).
Article 10(2)(g) AIA appears to be inspired by the technical literature on fairness metrics for algorithmic outputs. Nevertheless, Article 10(2)(g) AIA targets the input of algorithms. Therefore, classical fairness metrics for algorithmic outputs do not apply directly. On the other hand, input bias data detection, prevention, and mitigation methods remain largely unexplored. These methods are not clearly defined in the AI Act (Wachter 2024; Deck et al. 2024b).
Even if we assume that Article 10(2)(g) AIA is applicable to the output of an algorithm, the interpretation of its meaning is not self-evident. Although the wording in Article 10(2)(g) AIA does not directly relate to fairness metrics, from a teleological interpretation of the law, one might infer that Article 10(2)(g) AIA targets traditional output-related machine learning fairness metrics. In such cases, providers must ensure that the chosen metrics are suitable for detecting possible biases. As discussed above, there are many different ways to measure fairness in machine learning systems (Mehrabi et al. 2021). Therefore, it is already difficult to select a particular method to measure fairness. Furthermore, challenges arise beyond the choice of a specific measure. Even within a single measure, trade-offs may exist between the appropriateness of fairness metrics and other requirements such as accuracy (Mittelstadt, Wachter, and Russell 2023). For example, algorithms often show an inherent trade-off between accuracy and fairness in particular settings (see Figure 1). Although this trade-off can be reduced with approaches like “Levelling Up” trade-offs are difficult to navigate. Should developers focus more on accuracy (see also Article 15 AIA) or on the non-discrimination matters? Balancing these needs is a non-trivial task. Even for computer science experts, it is currently not entirely clear how the application of fairness metrics should be managed in the context of such trade-offs. Accordingly, the question persists: What constitutes an “appropriate measure” within the meaning of the AIA? The AIA’s legislative text is sufficiently indeterminate to permit varying interpretations concerning the precise objective and intended scope of the requirement for appropriate measures.
Figure 1: Algorithmic fairness vs. accuracy trade-off, data adopted from (Mittelstadt, Wachter, and Russell 2023). While the accuracy improves, the measure of algorithmic fairness (here demographic parity for the sex attribute) increases as well. In order to reduce algorithmic unfairness, the accuracy of the algorithm decreases. Permission to replot the data was obtained from the original authors.
However, compared to output fairness metrics, technical bias detection methods for input data are understudied. Furthermore, as previously noted, the causes of biases in ML are multi-causal (Gebru 2019; Gebru et al. 2021; Hanna et al. 2020; Mehrabi et al. 2021; Bender et al. 2021). Even when data are collected representatively, biases can occur (see (Mehrabi et al. 2021) for a comprehensive list). Two examples illustrate this argument:
Social bias can occur regardless of sample representativeness (Mitchell et al. 2021). For instance, in loan application settings, if certain groups (such as women) systematically receive lower wages, the data will contain inherent bias even with perfect representation.
Bias can emerge when one subgroup is smaller than others (Barocas, Hardt, and Narayanan 2023). Intuitively, an algorithm’s performance decreases when trained on fewer data points, resulting in poorer performance for smaller subgroups. Since smaller subgroups often correspond to minorities, algorithms may perform worse for these populations even with representative data.
The data governance strategies in Article 10(2)(g) AIA focus on the technical aspects of input data biases. While systematic fairness approaches have been developed for output bias detection, prevention, and mitigation (Han, Baldwin, and Cohn 2023; Krasanakis and Papadopoulos 2024; Pagano et al. 2023; Zollo et al. 2024), we lack similarly comprehensive approaches for input data analysis. This creates a significant gap between the legal requirements in Article 10(2)(g) AIA and their technical and ethical feasibility.
Article 10(2) AIA: Additional obligations without specific fairness regulations In reviewing of Article 10(2) AIA, it is noticeable that the other sections of Article 10(2) AIA also indirectly link to algorithmic fairness considerations. For example, the representativeness of training data is explicitly mentioned in Article 10(2)(d). When interpreting Article 10(2)(f) AIA or Article 10(2)(g) AIA, however, the additional requirements in Article 10(2) AIA do not result in any new insights. Since these articles have no direct requirements regarding non-discrimination, we will not analyze them in further detail.
Article 10(5) AIA: A legal basis for the processing of personal data in non-discrimination contexts. Finally, Article 10(5) AIA allows the processing of Article 9 GDPR data (see also Recital 70 AIA). Article 9 GDPR protects special categories of personal data, such as genetic, biometric, or health data (Article 9(1) GDPR). There is tension (Deck et al. 2024a) between the need for debiasing AI algorithms and data protection law, which Article 10(5) aims to address. In order to effectively mitigate biases in AI systems, the processing of personal data (for example, to compute fairness metrics) is important (van Bekkum 2024). Notable here is that this exception only applies in the high-risk regime. Thus, this exception is not applicable to non-high-risk systems. Nevertheless, Article 10(5) is an example of how the AI Act not only imposes burdens on developers and deployers of AI systems but also grants rights to these groups. This aspect is frequently underappreciated in the discourse regarding the AI Act.
Article 11 (1) AIA and Annex IV (2)(g) AIA: Technical documentation of bias testing methods
Article 11 (1) together with Annex IV (2)(g) AIA requires that the technical documentation of a high-risk system includes “information about the validation and testing data used and their main characteristics; metrics used to measure accuracy, robustness and compliance with other relevant requirements set out in Chapter III, Section 2, as well as potentially discriminatory impacts”.
From a grammatical interpretation, this section only specifies that the methods enforced in Chapter III, Section 2 AIA are documented. No new additional requirements are enforced. Would a teleological interpretation change this analysis? Even if Article 11(1) AIA, together with Annex IV (2)(g) AIA, were to enforce new measures, the requirement would again only target validation and testing data  [^13] In contrast to Article 10, AIA training data are not mentioned. . Thus, no additional and new requirements for high-risk AI systems on the output side are enforced through Article 11(1) AIA.
Article 13 AIA: Transparency and Provision of Information to Deployers Article 13(1) AIA provides that ‘High-risk AI systems shall be designed and developed in such a way as to ensure that their operation is sufficiently transparent to enable deployers to interpret a system’s output and use it appropriately”. One could contend that the interpretation of a system’s output necessarily entails an assessment of potential discrimination. Recital 72 enumerates aspects regarding the information that should be provided, with specific reference to the protection of fundamental rights. However, the risk identified in the recital is presented in general terms and fails to specify any particular methodology or technological tool that must be employed. Moreover, as Recitals serve a non-binding role and do not possess direct normative force in the interpretation of Article 13, the lack of specificity in Recital 72 does not alter the conclusion. Accordingly, Article 13(1) itself does not impose an obligation to use algorithmic fairness measures or metrics derived from the computer science domain.
Furthermore, even assuming, arguendo, that the use of such algorithmic fairness tools is required during the development or deployment process, Article 13(1) merely obliges providers to document the results. The imposition of mitigation or prevention measures does not fall within the mandatory scope of Article 13(1).
Article 15 (4) AIA: Output bias regulation in case of feedback loops. In contrast to Article 10(2)(f) AIA, Article 15(4) AIA mandates that “High-risk AI systems that continue to learn after being placed on the market or put into service shall be developed in such a way as to eliminate or reduce as far as possible the risk of possibly biased outputs influencing input for future operations (feedback loops) and as to ensure that any such feedback loops are duly addressed with appropriate mitigation measures.” In this case, the output of a system is regulated, in contrast to the input of an AI system in Article 10 AIA. The reasoning behind this is that AI systems should not become echo chambers for biases (Zuiderveen Borgesius et al. 2024). According to Recital 67 AIA, this is a special concern when examining historical biases. Most important to note here is the fact, that this output regulation only applies if the systems continue learning after being placed on the market or put into service.
Additionally, it remains unclear if the word “bias” refers to the same bias as in Article 10 AIA. Recital 67 AIA, which relates to Article 10 AIA but also mentions feedback loops, indicates that the same biases are addressed.
Biased output is usually detected and mitigated with fairness metrics. As shown above, the interpretation of fairness metrics is challenging as discussed above. Therefore, it becomes very challenging to identify how to “reduce as far as possible the risk of possibly biased outputs” in Article 15(4) AIA. This requirement is vague.
Furthermore, especially for language (generative) models, biases in the input data and biases in the output data might be different. In this context, the mention of bias in Article 15, titled “Accuracy, robustness, and cybersecurity”, is rather odd since the latter two fields are traditionally not mainly concerned with discrimination (Nolte, Rateike, and Finck 2024). Here, it remains an open question whether the robustness of an AI system and fairness represent a trade-off (Lee et al. 2021).
Other articles regarding high-risk systems regulations. While the previously discussed articles provide at least some intuition about addressed concerns, we categorize the remaining articles of Table 1. Most of these paragraphs only concern fundamental rights impacts, which show a limited impact on non-discrimination regulations. Thus, the discussion here is rather short. Article 13(3)(b)(iii) AIA mandates that information on the foreseeable misuse and the impact on fundamental rights is provided. Article 14(2) AIA mandates that human oversight is needed to minimize risks to fundamental rights. Article 17(1)(i) AIA and Article 26(5) AIA are related to reporting and information on serious incidents. Serious incidents are linked to discrimination through Article 3(49)(c) AIA: “’serious incident’ means an incident or malfunctioning of an AI system that directly or indirectly leads to any of the following: […] the infringement of obligations under Union law intended to protect fundamental rights”. Finally, Article 27(1) AIA requires a fundamental rights impact assessment when the high-risk system is used in special systems within bodies governed by public law. Yet again, significant gaps exist between theoretical and methodological elaboration of these impact assessments (Mantelero 2024). Article 27(1) AIA only addresses public parties. A fundamental right impact assessment is not necessary for private parties.
Finally, the impact of fundamental rights on non-discrimination is also monitored by a specific authority established through Article 77 (1) and Article 77 (2) AIA. The effect and impact of this authority framework should be assessed and evaluated after it is established. Summary: High-risk systems balancing between specificity and generality The AI Act aims to be a specific law for AI (Ződi 2024). For the ML and Law community, it might be challenging to understand the implications of this specific AI Law. Some Articles mention non-discrimination directly, while others focus on fundamental rights. Only Article 10 AIA and Article 15 AIA explicitly mention non-discrimination and biases, compared to the statement of fundamental rights protection in Article 9 AIA. However, Article 10 AIA only considers the input phase, while Article 15 AIA only addresses the output phase for feedback loops. It remains open why different Articles in the AI Act do not regulate the same types of direct, indirect, and intersectional discrimination on both the input and output sides.
Interplay between traditional EU non-discrimination law and the AI Act for high-risk systems
The AI Act focuses on the input side of high-risk systems. The output side is not covered in comparable detail. From a computer science perspective, research focuses on the output side of algorithms. As a lex specialis (Craig and De Búrca 2011), the AI Act prevails over general frameworks to the extent that it covers an issue; where it does not address certain discriminatory outcomes, however, the more general non-discrimination law remains applicable. Since output regulation is not present, traditional EU non-discrimination law could be applied.
The interplay between the AI regulations and fundamental rights requires further research, including clarification regarding regulation when the AI Act protects specific fundamental rights such as non-discrimination (Lewis et al. 2025). This also includes an analysis of when fundamental rights are vertically and horizontally applicable. Furthermore, the fundamental rights in the AI Act need to be balanced (Kusche 2024), and the material scope is limited due to the enumeration of protected attributes (Xenidis and Senden 2020). As pointed out, “gaps in the regulatory framework have left fundamental rights inconsistently protected in the AI Act.” (Palmiotto 2025) Nevertheless, since the AI Act does not protect individuals regarding the output of algorithms, there is room for the applicability of the CFR. It remains an open research question how the protection in the CFR and the AI Act relate (Lewis et al. 2025).
Besides the CFR, the protection of individuals from non-discrimination has been concretized in secondary EU law (Xenidis and Senden 2020; Lewis et al. 2025). Compared to the CFR, the horizontal applicability is clear. For example, Directives 2004/113/EC (Equal Treatment in Goods and Services Directive) and 2000/43/EC (Race and Ethnicity Equality Directive) specifically target non-discrimination regulations. These regulations address specific cases in which discrimination can occur. Due to their specificity, they have a limited material and personal scope as a result of the enumeration of protected attributes (Xenidis and Senden 2020; Cîrciumaru 2024). . It therefore remains to be seen how the interplay between the existing regulations, which cannot simply be transferred to AI, and the AI Act will develop. It also remains to be seen how national and EU courts will position themselves. In the past, the ECJ has already issued important rulings on how the non-discrimination provisions in Article 21 CFR must be interpreted in relation to secondary EU law (Xenidis and Senden 2020). It is very likely that in the future, the ECJ will continue to issue rulings to clarify the interplay between the EU AI Act, fundamental rights, and other directives and regulations. There is still a long way to go.
The AI Act holds the potential to complement traditional non-discrimination legislation. Traditional non-discrimination law primarily targets the protection of individuals. Despite the identified challenges associated with this individual-focused approach, it nonetheless provides a degree of protection. However, such an emphasis on individuals can overlook broader regulatory dimensions. In the context of EU data protection, it has been posited that a dual approach involving both individual protection and collaborative governance is necessary (Kaminski 2018). A dual approach might also be beneficial for non-discrimination regulations in the context of AI. The EU AI Act is fundamentally aligned with product safety regulation rather than individual rights protection (Almada and Petit 2023). The strength of this regulatory framework lies in its capacity to extend beyond individual-focused concerns. To maximize its efficacy, we contend that it is imperative to clarify the relationship between this framework and traditional non-discrimination regulations. It is plausible that both the product safety focus and individual rights considerations will be key requirements for the effective and comprehensive regulation of AI systems.
4 Summary & Outlook
As Mayson (2018) already pointed out: Bias in, bias out. While regulating these biases, AI encounters significant complexities.
First, the AI Act focuses on discrimination resulting from data. The non-discrimination regulations analyzed in the above section mostly focus on bias in (training) data and thus exclude other sources of bias, such as design choices in the algorithms. Furthermore, the regulations lack specificity in detection and mitigation methods. That makes it hard for practitioners to follow the law.
Second, the AI Act does not include many regulations for systems ranging from non-risk to limited-risk. The AI Act not only imposes obligations on high-risk systems but also grants rights, such as the authorization to use special categories of personal data (Article 9 GDPR) in Article 10(5) AIA in high-risk AI contexts. Such permission could also be granted for non-high-risk systems. Non-model specific regulations, such as the obligation to foster AI literacy in Article 4 AIA, are insufficient to ensure non-discriminatory practices. From our perspective, legislators missed an opportunity to regulate AI systems on a broader scope.
Many of the rules focus on (internal) compliance (Zuiderveen Borgesius et al. 2024) and checklists rather than on output-based safeguards for users affected by the AI system. This internal compliance is also linked to the self-assessment approach. It is questionable whether the self-assessment is suitable for protecting individuals and their fundamental rights as well as non-discrimination regulation (Bosoer, Cantero Gamito, and Rubio-Marin 2023).
Future. While the AI Act is casuistic and incomplete at the same time (Ződi 2024), we believe that the AI Act would benefit from four refinements. (1.) If the legislators decide to target non-discrimination regulation, non-discrimination should be explicitly mentioned. In contrast to terms such as bias or fairness, non-discrimination is a well-established legal concept within other European legislative frameworks. (2.) For a more consistent interpretation, it would be helpful if non-discrimination regulations cover input as well as output parts of the algorithms in one single article, also along the full AI life cycle (Deck et al. 2024b). Clarification is needed on appropriate measures for non-discrimination. Algorithmic output in terms of non-discrimination is only regulated if systems continuously learn after being placed on the market. Otherwise, outputs in terms of discriminatory results of AI systems are not regulated at all. Technical and non-technical means for bias detection, prevention, and mitigation for input and output data should be developed. For individuals, it could also be important to have a reverse burden-of-proof approach (Wachter 2024), making it easier to claim rights. (3.) Classical EU non-discrimination law will still be important in the AI context. More research is needed on how product safety-related regulations in the AI Act can be complied with by classical individual-based non-discrimination law. (4) Similar to the Code of Practice for GPAI models, it might be possible to start such a process also for high-risk systems. Technical and legal experts, as well as all stakeholders, could work on refined non-discrimination measures for high-risk systems. It is especially important to note that this work is best accomplished through joint community efforts between legal scholars and ML researchers. A bridge between these two fields is highly desirable since regulation should be guided by technical appropriateness and feasibility.
In the introduction of our work, we raised the question of how fair results of AI algorithms can be ensured. While the AI Act provides some direction, we also encounter significant challenges at the intersection of algorithmic fairness and non-discrimination law. Our paper aims to bridge these gaps between regulatory and technical limitations in the machine learning domain. From our perspective, both disciplines can benefit from closer interactions.
5 Acknowledgments
We would like to thank Miriam Rateike for her detailed comments on an earlier version of the manuscript. We are grateful for comments from Christoph Kern and the Social Data Science and Statistical Learning group during an institute seminar, which clearly improved this paper. Furthermore, we thank Alina Wenick, Tharushi Abeynayaka, and Mara Seyfert for their comments on the manuscript and for proofreading the paper.
Kristof Meding was supported by the Carl Zeiss Foundation through the CZS Center for AI and Law. Additionally, Kristof Meding is a member of the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645.